Shaniqua brings home 225$ per week working 25 hours. She is able to save 25$ of her earnings each week. Shaniqua wants to have 600$ at the end of 15 weeks. She is allowed to work up to 30 hours per week if she wants. She can save all the money earned working extra hours. Hoe many weeks will it take for Shaniqua to meet her goal
1
Expert's answer
2018-11-13T11:40:10-0500

If Shaniqua brings home 225$ per week working 25 hours, then her wage rate is 225/25 = $9 per hour.

If Shaniqua is allowed to work up to 30 hours per week if she wants and can save all the money earned working extra hours,

then she will be able to save 25 + 9*5 = $70 and she will need 600/70 = 8.57 or less then 9 weeks to meet her goal and have 600$.